JWars - A Generic Strategy Game in Java

Midterm Project at IMM

Michael Francker Christensen s031756 Ask Hjorth Larsen s021864

Supervisor: Paul Fischer DTU - Technical University of Denmark

10th May 2006

This space intentionally left blank 2

Table of contents

3